# Tailored Teaching with Balanced Difficulty: Elevating Reasoning in Multimodal Chain-of-Thought via Prompt Curriculum

### v1

@@ Abstract
Figure 1: The simplified flowchart of CAMS and its effects
Introduction
Multimodal Large Language Models (MLLMs)  (Yin et al. 2024; Liu et al. 2023) have emerged prominent research focus alongside the rapid advancement of artificial intelligence. Built upon powerful foundation language models  (Touvron et al. 2023; Bai et al. 2023), MLLMs leverage cross-modal alignment mechanisms to achieve understanding and processing of information across multiple modalities, including text, images, videos, and audio. A typical approach to deploying MLLMs is the in-context learning paradigm  (Brown et al. 2020; Xie and Min 2022), which drives models to perform predictions by providing a large number of instructions and input-output pair examples. Chain-of-Thought  (Wei et al. 2022; Zhou et al. 2022) enhances the logical reasoning capability of models by constructing examples that decompose complex problems into step-by-step subproblems and solve them sequentially. Multimodal Chain-of-Thought (MCoT)  (Zhang et al. 2023), on the other hand, further extends this core idea to the application scenarios of multimodal large language models, achieving an improvement in cross-modal reasoning capabilities.
MCoT encourages large language models to perform multi-step reasoning by providing explicit problem-solving rationales, rather than mapping questions directly to answers. This approach has been effectively applied to scientific question answering across domains such as natural sciences, linguistic sciences, and social sciences  (Lu et al. 2022; Chen et al. 2024; Goyal et al. 2017; Marino et al. 2019; Schwenk et al. 2022). However, as illustrated in Figure 2(a), MCoT prompt examples are typically randomly selected or manually crafted without considering the model’s internal knowledge distribution or the characteristics of the dataset. This leads to prompts that are unstable and insufficiently tailored to the model or task. As a result, MCoT prompting often suffers from poor generalization, hallucinated outputs, and inconsistent performance. To enhance the specificity and effectiveness of prompt examples, Auto-CoT  (Zhang et al. 2022) constructs prompts by selecting representative examples via clustering, as shown in Figure 2(b). However, due to the inherent differences between modalities, the effectiveness of Auto-CoT in multimodal scenarios is limited. It fails to adequately account for the distributional differences in internal knowledge across different models.
@@ Introduction
We propose CAMS (Complexity-Guided Active Multimodal CoT Sampling), a novel joint selection framework that integrates active learning and data complexity assessment. Rather than relying on randomly selected or manually crafted MCoT examples, CAMS constructs a tailored “error sets” for each model based on the training sets of various reasoning tasks, aiming to enhance the models’ performance on test sets. As shown in Figure 2(c), CAMS identifies highly targeted, effective, and difficulty-balanced prompt examples by jointly analyzing sample uncertainty and complexity. Figure 1(a) illustrates the core workflow of CAMS, while Figure 1(b) highlights its perforamance advantages. We conduct experiments on five benchmark datasets and three multimodal large models. The results show that CAMS both improves model performance and greatly reduces accuracy variability caused by random prompt selection. Our key contributions are as follows:
@@
Related Work
Prompt Selection based on Active Learning
Active learning is a machine learning paradigm focused on maximizing model performance using the fewest possible labeled samples. It aims to identify the most informative unlabeled data points for annotation, thereby reducing labeling costs while maintaining high accuracy. Active learning methods are typically grouped into the following three categories based on how unlabeled data is queried: membership query synthesis (Angluin 1988; King et al. 2004), where the model generates new instances for labeling; stream-based selective sampling (Dagan and Engelson 1995; Krishnamurthy 2002), where data points are evaluated one at a time for potential labeling; and pool-based sampling (Lewis 1995), where the model selects the most informative samples from a large pool of unlabeled data.
Active Prompt (Diao et al. 2023) applies active learning principles to prompt selection by quantifying uncertainty through prediction disagreement metrics (e.g., variance, entropy, disagreement) and choosing high-uncertainty samples as prompt examples. However, Active Prompt focuses solely on model’s prediction disagreement with samples (i.e., distributional differences in model-internal knowledge) without considering sample complexity. Our approach introduces a data complexity evaluator that assesses the inherent difficulty of samples. This allows for more customized and effective prompt selection, combining insights from both model knowledge and dataset characteristics.
Chain-of-Thought in Visual Question Answering
The Multimodal Chain of Thought (CoT) technique is widely adopted to enhance the multi-step reasoning abilities of large language models (LLMs). Its core idea is to guide models to generate intermediate reasoning steps that help them solve complex problems more effectively. Benchmarks such as VQA (Antol et al. 2015), VQAv2 (Goyal et al. 2017), OK-VQA (Marino et al. 2019), A-OKVQA (Schwenk et al. 2022), and ScienceQA (Lu et al. 2022) provide structured visual question answering (VQA) tasks across various domains, including natural science, social science, semantics, and everyday reasoning. For complex reasoning tasks, recent approaches like MCoT (Zhang et al. 2023), Auto-CoT (Zhang et al. 2022), Self-Consistency (Wang et al. 2022), and Active Prompt (Diao et al. 2023) leverage carefully designed prompt examples to improve model performance.
Despite these advancements, many methods (Wei et al. 2022; Wang et al. 2022; Zhou et al. 2022) rely on either randomly selected or manually crafted prompt examples. These examples often fail to align with the specific demands of individual VQA tasks, limiting model performance. In particular, they tend to overlook key factors such as the distributional characteristics of the model’s internal knowledge and the multimodal nature of VQA tasks, focusing primarily on unimodal scenarios. To address these limitations, our framework jointly considers the model’s uncertainty about the dataset and the intrinsic complexity of each example. By integrating both model-centric and data centric perspectives, we dynamically construct prompt examples that are better tailored to the model’s current capabilities and the reasoning requirements of the task, leading to more effective and adaptive prompting in visual question answering.
Methodology
Figure 3 illustrates the three core modules of CAMS: (i) Analysis of Multimodal Model Internal Knowledge, which quantifies the model’s predictive uncertainty through multiple independent samplings of the model, revealing the distribution characteristics of the model’s internal knowledge; (ii) Complexity-Based Dataset Feature Estimation, which is used to quantitatively evaluate dataset characteristics; (iii) Examples Sampling Strategy, which incorporates model uncertainty indicators, dataset complexity scores, and the “easy-hard” selection principle to dynamically identify the most representative and effective prompt examples for the target model.
Problem Definition
@@
where $\mathcal{P}_{space}$ denotes the set of all possible prompts (automatically selected or manually crafted), and $\mathcal{S}$ represents the corresponding evaluation metric.
Analysis of Multimodal Model Internal Knowledge
Disagrement.
@@
where $a_{i}=\{\mathcal{F}^{1}_{\theta}(q_{i}),\mathcal{F}^{2}_{\theta}(q_{i}),...,\mathcal{F}^{k}_{\theta}(q_{i})\}$ represents the k sampling outcomes for sample $q_{i}$. We then compute unique answer via set operations to remove duplicates, yielding $k^{\prime}$ unique items $a_{i}=\{\mathcal{F}^{1}_{\theta}(q_{i}),\mathcal{F}^{2}_{\theta}(q_{i}),...,\mathcal{F}^{k^{\prime}}_{\theta}(q_{i})\}$, where $k^{\prime}$ denotes the disagrement metric $u$
Clustering.
After assigning each training sample a corresponding uncertainty metric $u$, we perform clustering on all training samples based on the numerical values of $u$, yielding a new training sample set:
$\mathcal{D}^{n}_{train}=\{\mathcal{D}_{u_{1}},\mathcal{D}_{u_{2}},...,\mathcal{D}_{u_{n}}\}$ (4)
where $n$ denotes the total number of distinct uncertainty metrics $u$ in the training samples, $u_{i}$ represents the specific uncertainty metric (Disagrement) satisfying the order $u_{1}>u_{2}>...>u_{n}$, and $\mathcal{D}_{u_{i}}$denotes the set of all training samples with $Disagrement=u_{i}$. To further facilitate difficulty grading of training samples, we partition the samples according to their uncertainty levels. The first $n/2$ sets are classified as difficult error-prone questions, while the subsequent $n/2$ sets are defined as simple fundamental questions, resulting in a difficulty-divided dataset:
@@
$\mathcal{D}_{difficulty}$ $=\{\mathcal{D}_{u_{1}},\mathcal{D}_{u_{2}},\dots,\mathcal{D}_{u_{\lfloor n/2\rfloor}}\},$
$\mathcal{D}_{easy}$ $=\{\mathcal{D}_{u_{\lfloor n/2\rfloor+1}},\dots,\mathcal{D}_{u_{n}}\}.$
@@
Complexity-Based Dataset Feature Estimation
Evol Complexity.
This is an evolution-based metric. We define a small-scale seed dataset $\mathcal{D}=\{(I^{(0)}_{1},R^{(0)}_{1}),(I^{(0)}_{2},R^{(0)}_{2}),...,(I^{(0)}_{N},R^{(0)}_{N})\}$, where $(I^{(0)}_{i},R^{(0)}_{i})$ denotes instruction-response pairs. For each instruction sample $I$ , we enhance its complexity through Technique $\mathcal{F}_{\alpha}(\cdot)$, which involves adding constraints, specification, and increasing reasoning steps. After $N$ iterations, a set of instructions with varying complexities is obtained:
where $I^{(m)}_{i}=\mathcal{F}_{\alpha}(I^{(m-1)}_{i})$ and $N$ is set to 5. Further, we utilize the scoring function $\mathcal{S}(\cdot)$ (i.e. ChatGPT) to rate and rank these 6 samples, generating a set of instructions with scoring labels:
$\{(I^{(j)}_{i},\mathcal{S}(I^{(j)}_{i}),R^{(j)}_{i})\}^{M}_{j=0}$ (7)
Composed of these labeled instruction groups, a dataset is constructed to train llava as the final complexity scorer.
Textual Substitution for Visual Content.
This scheme aims to convert images in multimodal data into text form (image captions) and integrate them with question texts to form unimodal inputs, enabling the reuse of existing unimodal complexity scorers. Specifically, we first employ the ViT model (Dosovitskiy et al. 2020) to generate a corresponding text description (caption) for each image in the multimodal dataset. The image captions are then integrated with the corresponding questions and options to form complete unimodal inputs in pure text. We adopt the concatenation format of “question text + option text + image caption”, using the delimiter ${\prime}\backslash n^{\prime}$ to ensure that the scorer can recognize the text representation of visual information. Finally, the integrated text data is fed into the complexity scorer, which outputs the corresponding sample complexity score.
Strategies for Selecting Examples
We employ a difficulty-balanced sample selection strategy. Specifically, for subsets of questions categorized as difficult and error-prone, we select an equal number of high-complexity and low-complexity questions; this balanced selection criterion is equally applied to subsets of simple and fundamental questions. To elaborate, we take $n=4$ as example. Let the uncertainty metrics be $u_{1}=4,u_{2}=3,u_{3}=2,u_{4}=1$. The training sample set after categorization can be formalized as $\mathcal{D}^{n}_{train}=\{\mathcal{D}_{4},\mathcal{D}_{3},\mathcal{D}_{2},\mathcal{D}_{1}\}$, where $\mathcal{D}_{4}$ and $\mathcal{D}_{3}$ are defined as difficult and error-prone question subsets owing to their high uncertainty metrics, while $\mathcal{D}_{2}$ and $\mathcal{D}_{1}$ are classified as simple and fundamental question subsets due to their low uncertainty metrics. When selecting four training samples as prompt examples, one feasible strategy is to choose questions with the highest complexity scores from $\mathcal{D}_{4}$ and $\mathcal{D}_{2}$, paired with questions with the lowest complexity scores from $\mathcal{D}_{3}$ and $\mathcal{D}_{1}$. Alternatively, an inverse selection scheme can be adopted: selecting questions with the highest complexity scores from $\mathcal{D}_{3}$ and $\mathcal{D}_{1}$, and those with the lowest complexity scores from $\mathcal{D}_{4}$ and $\mathcal{D}_{2}$.
Experiments
  1. Q1.
  2. Q2.
  3. Q3.
  4. Q4.
  5. Q5.
Experimental Settings
Datasets and Metrics.
Following the standard evaluation settings in MLLMs reasoning research,we conduct testing on five popular benchmarks: ScienceQA (Lu et al. 2022), A-OKVQA (Schwenk et al. 2022), OK-VQA (Marino et al. 2019), VQAv2 (Goyal et al. 2017) and TextVQA (Singh et al. 2019). The problem designs in these datasets include both multiple-choice and open-ended answer formats, and we follow the practice (Liu et al. 2024; Li et al. 2024; Bai et al. 2023) of using accuracy as the metric.

### v2

@@ Abstract
1 Introduction
Multimodal Large Language Models (MLLMs)  Yin et al. (2024); Liu et al. (2023a) have emerged prominent research focus alongside the rapid advancement of artificial intelligence. Built upon powerful foundation language models  Touvron et al. (2023); Bai et al. (2023), MLLMs leverage cross-modal alignment mechanisms to achieve understanding and processing of information across multiple modalities, including text, images, videos, and audio. A typical approach to deploying MLLMs is the in-context learning paradigm  Brown et al. (2020); Xie & Min (2022), which drives models to perform predictions by providing a large number of instructions and input-output pair examples. Chain-of-Thought  Wei et al. (2022); Zhou et al. (2022) enhances the logical reasoning capability of models by constructing examples that decompose complex problems into step-by-step subproblems and solve them sequentially. Multimodal Chain-of-Thought (MCoT)  Zhang et al. (2023), on the other hand, further extends this core idea to the application scenarios of multimodal large language models, achieving an improvement in cross-modal reasoning capabilities.
MCoT encourages large language models to perform multi-step reasoning by providing explicit problem-solving rationales, rather than mapping questions directly to answers. This approach has been effectively applied to scientific question answering across domains such as natural sciences, linguistic sciences, and social sciences. However, as illustrated in Figure 1(a), MCoT prompt examples are typically randomly selected or manually crafted without considering the model’s internal knowledge distribution or the characteristics of the dataset. This leads to prompts that are unstable and insufficiently tailored to the model or task. As a result, MCoT prompting often suffers from poor generalization, hallucinated outputs, and inconsistent performance. To enhance the specificity and effectiveness of prompt examples, Auto-CoT  Zhang et al. (2022) constructs prompts by selecting representative examples via clustering, as shown in Figure 1(b). However, due to the inherent differences between modalities, the effectiveness of Auto-CoT in multimodal scenarios is limited. It fails to adequately account for the distributional differences in internal knowledge across different models.
@@ 1 Introduction
We propose CAMS (Complexity-Guided Active Multimodal CoT Sampling), a novel joint selection framework that integrates active learning and data complexity assessment. Rather than relying on randomly selected or manually crafted MCoT examples, CAMS constructs a tailored “error sets” for each model based on the training sets of various reasoning tasks, aiming to enhance the models’ performance on test sets. As shown in Figure 1(c), CAMS identifies highly targeted, effective, and difficulty-balanced prompt examples by jointly analyzing sample uncertainty and complexity. We conduct experiments on five benchmark datasets and three multimodal large models. The results show that CAMS both improves model performance and greatly reduces accuracy variability caused by random prompt selection. Our key contributions are as follows:
@@
2 Related Work
2.1 Prompt Selection based on Active Learning
Active learning is a machine learning paradigm focused on maximizing model performance using the fewest possible labeled samples. It aims to identify the most informative unlabeled data points for annotation, thereby reducing labeling costs while maintaining high accuracy. Active learning methods are typically grouped into the following three categories based on how unlabeled data is queried: membership query synthesis Angluin (1988); King et al. (2004), where the model generates new instances for labeling; stream-based selective sampling Dagan & Engelson (1995); Krishnamurthy (2002), where data points are evaluated one at a time for potential labeling; and pool-based sampling Lewis (1995), where the model selects the most informative samples from a large pool of unlabeled data.
Active Prompt Diao et al. (2023) applies active learning principles to prompt selection by quantifying uncertainty through prediction disagreement metrics (e.g., variance, entropy, disagreement) and choosing high-uncertainty samples as prompt examples. However, Active Prompt focuses solely on model’s prediction disagreement with samples (i.e., distributional differences in model-internal knowledge) without considering sample complexity. Our approach introduces a data complexity evaluator that assesses the inherent difficulty of samples. This allows for more customized and effective prompt selection, combining insights from both model knowledge and dataset characteristics.
2.2 Chain-of-Thought in Visual Question Answering
The Multimodal Chain of Thought (CoT) technique is widely adopted to enhance the multi-step reasoning abilities of large language models (LLMs). Its core idea is to guide models to generate intermediate reasoning steps that help them solve complex problems more effectively. Benchmarks such as VQA Antol et al. (2015), VQAv2 Goyal et al. (2017), OK-VQA Marino et al. (2019), A-OKVQA Schwenk et al. (2022), and ScienceQA Lu et al. (2022) provide structured visual question answering (VQA) tasks across various domains, including natural science, social science, semantics, and everyday reasoning. For complex reasoning tasks, recent approaches like MCoT Zhang et al. (2023), Auto-CoT Zhang et al. (2022), Self-Consistency Wang et al. (2022), and Active Prompt Diao et al. (2023) leverage carefully designed prompt examples to improve model performance.
Despite these advancements, many methods Wei et al. (2022); Wang et al. (2022); Zhou et al. (2022) rely on either randomly selected or manually crafted prompt examples. These examples often fail to align with the specific demands of individual VQA tasks, limiting model performance. In particular, they tend to overlook key factors such as the distributional characteristics of the model’s internal knowledge and the multimodal nature of VQA tasks, focusing primarily on unimodal scenarios. To address these limitations, our framework jointly considers the model’s uncertainty about the dataset and the intrinsic complexity of each example. By integrating both model-centric and data centric perspectives, we dynamically construct prompt examples that are better tailored to the model’s current capabilities and the reasoning requirements of the task, leading to more effective and adaptive prompting in visual question answering.
@@ 2.2 Chain-of-Thought in Visual Question Answering
3 Methodology
Figure 2 illustrates the three core modules of CAMS: (i) Characterization of Multimodal Model Internal Knowledge, which quantifies the model’s predictive uncertainty through multiple independent samplings of the model, revealing the distribution characteristics of the model’s internal knowledge; (ii) Complexity-Based Dataset Feature Estimation, which is used to quantitatively evaluate dataset characteristics; (iii) Examples Sampling Strategy, which incorporates model uncertainty indicators, dataset complexity scores, and the “easy-hard” selection principle to dynamically identify the most representative and effective prompt examples for the target model.
3.1 Problem Definition
@@
where $\mathcal{P}_{space}$ denotes the set of all possible prompts (automatically selected or manually crafted), and $\mathcal{S}(\cdot)$ represents the corresponding evaluation metric.
3.2 Characterization of Multimodal Model Internal Knowledge
3.2.1 Disagreement.
@@ 3.2 Characterization of Multimodal Model Internal Knowledge
where $a_{i}=\{\mathcal{F}^{1}_{\theta}(q_{i}),\mathcal{F}^{2}_{\theta}(q_{i}),...,\mathcal{F}^{k}_{\theta}(q_{i})\}$ represents the k sampling outcomes for sample $q_{i}$, $\mathcal{F}^{j}_{\theta}(q_{i})$ denotes the j-th sample generated by the model for input, and $\mathcal{A}$ represents the collection of $k$ sampling results for all samples. We then compute unique answer via set operations to remove duplicates, yielding $k^{\prime}$ unique items $a_{i}=\{\mathcal{F}^{1}_{\theta}(q_{i}),\mathcal{F}^{2}_{\theta}(q_{i}),...,\mathcal{F}^{k^{\prime}}_{\theta}(q_{i})\}$, where $\frac{k^{\prime}}{k}\in(0,1.0]$ denotes the disagreement metric $u$. A larger value of the disagreement metric $u$ indicates a higher difficulty level for the sample.
3.2.2 Clustering.
After assigning each training sample a corresponding uncertainty metric $u$, we perform clustering on all training samples based on the numerical values of $u$, yielding a new training sample set. Specifically, we classify samples with $u\in(0.5,1.0]$ as difficult questions; conversely, samples with $u\in(0,0.5]$ are classified as easy questions.
@@
$\mathcal{D}_{difficulty}$ $=\{\mathcal{D}_{u_{i}}\}_{0.5<u_{i}\leq 1.0},$
$\mathcal{D}_{easy}$ $=\{\mathcal{D}_{u_{j}}\}_{0<u_{j}\leq 0.5}.$
where $u_{i}$ represents the specific disagreement value, and $\mathcal{D}_{u_{i}}$ represents the set of samples whose disagreement metric equals $u_{i}$.
3.3 Complexity-Based Dataset Feature Estimation
Training of the Complexity Scorer.
We introduce an evolution-based metric, Evol Complexity. We define a small-scale seed dataset $\mathcal{D}=\{(I^{(0)}_{1},R^{(0)}_{1}),(I^{(0)}_{2},R^{(0)}_{2}),...,(I^{(0)}_{N},R^{(0)}_{N})\}$, where $(I^{(0)}_{i},R^{(0)}_{i})$ denotes instruction-response pairs and $N$ denotes the number of instruction-response pairs. For each instruction sample $I$ , we enhance its complexity through Technique $\mathcal{F}_{\alpha}(\cdot)$, which involves adding constraints, specification, and increasing reasoning steps. After $N$ iterations, a set of instructions with varying complexities is obtained:
where $I^{(m)}_{i}=\mathcal{F}_{\alpha}(I^{(m-1)}_{i})$ , $M$ denotes the number of sample evolution iterations and is set to 5. Further, we utilize the scoring function $\mathcal{S}(\cdot)$ (i.e. ChatGPT) to rate and rank these 6 samples, generating a set of instructions with scoring labels:
$\{(I^{(j)}_{i},\mathcal{S}(I^{(j)}_{i}),R^{(j)}_{i})\}^{M}_{j=0}$ (6)
Composed of these labeled instruction groups, a dataset is constructed to train LLaMA as the complexity scorer. We use the Llama2-7B model trained on the “6k SFT + 10k DPO” dataset as the final complexity scorer. Experiments demonstrate that this model can effectively classify sample complexity, with an AlpacaEval score of 90.06% Liu et al. (2023b).
Computation of Complexity.
This scheme aims to convert images in multimodal data into text form (image captions) and integrate them with question texts to form unimodal inputs, enabling the reuse of existing unimodal complexity scorers. Specifically, we observe that each image in every multimodal dataset is accompanied by a corresponding textual description (i.e., a caption), and we thus decide to make full use of these image captions. The image captions are then integrated with the corresponding questions and options to form complete unimodal inputs in pure text. We adopt the concatenation format of “question text + option text + image caption”, using the delimiter ${\prime}\backslash n^{\prime}$ to ensure that the scorer can recognize the text representation of visual information. Finally, the integrated text data is fed into the complexity scorer, which outputs the corresponding sample complexity score.
3.4 Strategies for Selecting Examples
We employ a difficulty-balanced sample selection strategy. Specifically, for subsets of questions categorized as difficult and error-prone, we select an equal number of high-complexity and low-complexity questions; this balanced selection criterion is equally applied to subsets of simple and fundamental questions. This selection strategy fully incorporates two dimensions: uncertainty and complexity. !!!! It is important to note that all test samples within a given dataset share the prompt examples selected through this method, rather than each test sample using a distinct set of prompt examples.
4 Experiments
  1. Q1.
  2. Q2.
  3. Q3.
  4. Q4.
  5. Q5.
4.1 Experimental Settings
4.1.1 Datasets and Metrics.
Following the standard evaluation settings in MLLMs reasoning research,we conduct testing on five popular benchmarks: ScienceQA Lu et al. (2022), A-OKVQA Schwenk et al. (2022), OK-VQA Marino et al. (2019), VQAv2 Goyal et al. (2017) and TextVQA Singh et al. (2019). The problem designs in these datasets include both multiple-choice and open-ended answer formats, and we follow the practice Liu et al. (2024); Li et al. (2024); Bai et al. (2023) of using accuracy as the metric.
4.1.2 Baselines and Model Variants.
In our experiment, the following five methods are used as the main baselines: Chain-of-thought (ZS-CoT) Wei et al. (2022), Few shot CoT (FS-CoT), Self-consistency (Self-Con) Wang et al. (2022), Auto-CoT Zhang et al. (2022) and Active Prompt (Active-Pro) Diao et al. (2023). FS-CoT uses the same annotation process as our method, the only difference being that it randomly selects questions from the training data for annotation. We select multiple cutting-edge MLLMs as experimental models, including llama3.2-vision:11b, llava:7b, and Qwen2.5-VL:7b.
4.2 Implementation Details
4.2.1 Hyperparameters.
In the training set sampling phase, we follow the optimal value of 10 Diao et al. (2023) for the number of sampling iterations. During both the sampling and the inference phases of the experiments, the temperature is uniformly set at $0.5$. Unless otherwise specified, the multimodal large language models used in the experiments are llama3.2-vision:11b, llava:7b, and Qwen2.5-VL:7b.
4.2.2 Uncertainty Assessment.
In the experimental process, we adopt a zero-shot sampling strategy, which does not rely on additional examples or guidance information. For ScienceQA, A-OKVQA, and VQA, we perform sampling on the complete training sample sets, while for the VQAv2 and TextVQA datasets, we only sample 10,000 samples [^1] Because the training sets of these two datasets are too large, full sampling would require substantial computational and time costs.. For the sampling frequency parameter, We follow the optimal value Diao et al. (2023) to set $k=10$. We consistently use “Disagreement” [^2] In preliminary experiments, variance and entropy exhibit suboptimal performance in multimodal scenarios.—a more intuitive and accurate method—as the uncertainty metric.
4.2.3 Constructing Examples.
We focus on the innovating and optimizating strategies for prompt example selection. To reduce human labor as much as possible, we eliminate the need for manual annotation of prompt examples. Instead, we construct prompt instances by directly concatenating the question, reasoning, and answer from the dataset. The solution processes in the dataset are carefully crafted and rigorously screened by the authors, considering domain knowledge or task-specific styles. This ensures the generalization ability of our method to a certain extent. Details are provided in Appendix A.1.
@@ 4.2 Implementation Details
4.3 Main Results
4.3.1 For Q1: CAMS consistently outperforms nearly all baseline methods.
4.3.2 For Q2: CAMS can greatly reduce accuracy instability caused by randomly selecting prompt examples.